# Attention Retrieves, MLP Memorizes: Disentangling Trainable Components in the Transformer

### v2

@@ Abstract
The Transformer architecture is central to the success of modern Large Language Models (LLMs), in part due to its surprising ability to perform a wide range of algorithmic tasks – including mathematical reasoning, memorization, and retrieval – using only gradient-based training on next-token prediction. While the core component of a Transformer is the self-attention mechanism, we question how much, and which aspects, of the performance gains can be attributed to it. To this end, we compare standard Transformers to variants in which either the multi-layer perceptron (MLP) layers or the attention projectors (queries and keys) are frozen at initialization. To further isolate the contribution of attention, we introduce MixiT – the Mixing Transformer – a simplified, principled model in which the attention scores are entirely random and fixed at initialization, eliminating any input-dependent computation or learning in attention. Surprisingly, we find that MixiT matches the performance of fully trained Transformers on various algorithmic tasks, especially those involving basic arithmetic or focusing heavily on memorization. For retrieval-based tasks, we observe that having input-dependent attention scores is consistently beneficial, while MixiT underperforms. We attribute this failure to its inability to form specialized circuits such as induction heads – a specific circuit known to be crucial for learning and exploiting repeating patterns in input sequences. Even more interestingly, we find that attention with frozen key and query projectors is not only able to form induction heads, but can also perform competitively on language modeling. Our results underscore the importance of architectural heterogeneity, where distinct components contribute complementary inductive biases crucial for solving different classes of tasks. [^1] Our code is publicly available at https://github.com/princeton-pli/MixiT.
@@ 1 Introduction
Transformers [50] have rapidly become the workhorse architecture in modern machine-learning systems, powering state-of-the-art models in language, vision, and scientific domains [15, 48, 22]. Their success is typically attributed to the self-attention mechanism, which allows every token to aggregate information from the entire sequence and has been linked to emergent abilities such as long-range retrieval, algorithmic reasoning, and in-context learning. Yet we lack a precise answer to a fundamental question: which degrees of freedom inside the Transformer are truly necessary for these behaviours, and which can be simplified away without harming performance?
Prior work has probed the internals of trained Transformers. Studies of attention maps consistently report the emergence of induction heads that copy information forward and enable in-context retrieval [37]. Another line of work has focused on whether and how Transformers’ capabilities can emerge [1, 41, 24] by studying synthetic datasets with Zhong and Andreas [59] showing that models with frozen self-attention but trainable embeddings still solve many algorithmic tasks. These results hint that different parts of the architecture are responsible for modeling different tasks.
A standard Transformer block, however, contains several interacting components: an attention block, with learnable and input-dependent queries, keys form the attention weights, and MLP blocks composed of fully connected layers. To exemplify the complexity of the interaction, notice that the attention weights can change their value both through changes in the queries/keys parameters, as well as as through the residual-stream representations that feed those projections. Hence, because all the components are trained together, it is hard to discern the contribution of each individual one to the solution of a given task. In this work, we study the role of each Transformer’s component by freezing different parts of the architecture to their value at initialization. In partiucular, we consider the following simplified variants:
@@
  • MixiT (Mixing Transformer), a Transformer variant in which the attention sub-layer is replaced by a fixed, randomly initialised mixing matrix. After initialisation the mixing weights are frozen, preventing the model from adapting interactions during training or conditioning them on input content. Learning therefore takes place exclusively in the embedding and MLP blocks and in residual branches.
@@
We compare across several categories of tasks: mathematical reasoning, sentiment classification, memorization, information retrieval and language modeling. This way, we can inspect whether and how the different parts of the architecture are able to solve primitive tasks that relate to a range of basic reasoning and memorization skills. Here we summarize our main contributions:
  • We present the surprising finding that Frozen-QK can perform competitively with the standard transformer on language modeling tasks. Indeed, Frozen-QK develops the ability to form specialized circuits such as induction heads during training.
@@
  • When tested on tasks that require forming induction heads, MixiT fails to achieve good performance, while both the fully trained transformer and Frozen-QK can solve them to close-to perfect accuracy.
  • On pure memorization tasks, Frozen-MLP loses two-thirds of the storage capacity of the fully-trainable Transformer, while Frozen-QK and MixiT only a third. This suggests that the MLP layers are largely responsible, but collaborates with attention, on memorization tasks.
Our results indicate that the Transformer architecture has some built-in inductive bias towards algorithmic abilities, namely the ability to form circuits. Furthermore, while for a variety of tasks no particular structure in the attention maps is needed, some learnable attention mechanism is required when flexibly copying tokens. As a consequence, due to the observed gap in language modeling, we speculate that the ability to solve associative recall and complex retrieval is a key ingredient to modern language models. More broadly, our results underpin the importance of having heterogeneous components that are more specialized in efficiently solving modular subtasks, enabling the model to express complex, wide-ranging behaviors.
@@ 2 Models Specification
We consider a decoder-only Transformer based on the widely used Llama architecture [49]. Given an input sequence $x\in\mathbb{R}^{V\times m}$, where $V$ is the vocabulary size and $m$ is the sequence length, embedded with a linear map to the hidden states $h_{0}=W_{emb}\,x$ where $W_{emb}\in\mathbb{R}^{V\times n}$, where $n$ is the width of the model. At its core, the transformer uses $L$ stacked modules alternating the causal multi-head self-attention layers and MLP layers. Each self attention head is defined as:
@@
where $W_{\ell}^{G},W_{\ell}^{U}\in\mathbb{R}^{{n_{m}\times n}}$ and $W^{D}_{\ell}\in\mathbb{R}^{{n\times n_{m}}}$. $n_{h}$ and $n_{m}$ are the dimensions of the queries/keys for each head and MLP projections. We also use skip connections in both blocks with a pre-normalization scheme [52] and causal masking. We apply rotary embeddings [45] to the queries and keys of each layer. As in common practice, we use $n_{m}=4n$ and $Hn_{h}=n$.
@@
We also design a model where the attention map itself is frozen and, to achieve that, input-independent. In the simplest case, this can be obtained by having a random matrix $M_{\ell}\in\mathbb{R}^{m\times m}$ entries with $\mathcal{N}(0,1/m)$ entries, where the factor of $1/m$ acts as a variance-preserving normalizer. To ensure a stable forward pass in terms of depth and width scaling, we follow the principles of attention shaping [35] and propose the following:
@@
where $W^{M}_{\ell}$ is frozen at initialization and $\bar{W}^{M}_{\ell}$ contains the column-wise empirical average of $W^{M}_{\ell}$, to ensure that each row sums up to 1. In the Appendix, , we show that this attention has a stable forward pass, in the sense that the kernel of the activations has a well-defined depth-and-width limit, converging to a stochastic differential equation (SDE) [29, 35]. When we adopt this architecture, all the weights, excluding the random attention matrix, are trainable. Notably, the following convergence result implies the stability of the forward pass, in particular ruling out the numerical degeneracy such as rank collapse and vanishing gradients [14, 34].
@@
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation 2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
@@
As the random attention matrix $I+\frac{1}{\sqrt{mn}}W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT does not depend on the input, the rotary positional embedding [45] standard in Llama models cannot be applied to MixiT, as rotary embeddings are added to learned key and query embeddings, which are not present in MixiT. Hence, we implement a learnable positional embedding for each token position in the sequence, and add it to the corresponding token embedding in the first layer.
@@ 3.1 Tasks
We benchmark on a variety of tasks, covering categories including basic mathematical reasoning, memorization, sentiment classification, and language modeling. The basic mathematical reasoning and memorization tasks are based on tasks used in [59], with increased difficulty on some tasks to better reflect differences between architectures. More experimental details on these tasks can be found in Appendix B.
@@
Each training instance encodes a small randomly generated sequence of pairs followed by a single query, and the model is required to emit the value associated with that query one step later. We uniformly sample a sequence length $m\sim\mathcal{U}\{1,\dots,m_{\max}\}$, where $m_{\max}$ is the maximum sequence length. We then sample $m$ keys $\{k_{i}\}_{i=1}^{m}$ iid from the set $\{\frac{V}{2},\dots,V\}$ and $m$ values $\{v_{i}\}_{i=1}^{m}$ from $\{1,\dots,\tfrac{\text{V}}{2}-1\}$, where $V=256$ is the vocabulary size. The resulting keys are interleaved with their values to form the prefix $[(k_{1},\,v_{1}),\,(k_{2},\,v_{2}),\,\dots,\,(k_{m},\,v_{m})]$. A query key $k_{q}$ with $q\in[m]$ is chosen uniformly at random from the keys, and appended to the sequence. The goal is to predict the value corresponding to the query token. This task isolates retrieval ability, and probes associative recall. We sample 40,000 sequences for training, and 4,000 for testing. Transformer-based models typically solve this task by forming induction heads [37].
@@
Following [42], the $k$-hop induction heads task, or hopk, recursively completes bigrams auto-regressively, by repeatedly predicting the token that followed the last occurrence of the currently-considered token. As an example: given the input $X=a{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}dc}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}ad}a$, the 2-hop induction heads prediction is ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c}$. [42] showed that hopk is solvable by a $O(\log k)$-depth transformer. Hence to achieve fair comparison across architectures, we fix the model depth at 5 layers, and search over other hyperparameters.
@@
[…]
@@
[…]
@@
We use the Yelp polarity reviews dataset [58] to test each model variation’s ability to predict a review’s sentiment, i.e. whether a review is positive or negative.
@@
To test the model’s ability to model natural language, we train the model to perform next-token-prediction. We use two datasets, Wikitext-103 [32] and Fineweb-edu [38]. Wikitext-103 consists of 1,801,350 cleaned-up Wikipedia articles, with test set of size 4358. And Fineweb-edu consists of top-quality entries collected from web-crawled data, focusing on educational text. We randomly sample 1,048,576 entries for training, and 4096 to test.
@@ 3.2 Model training
For each task, we perform a grid search over a range of hyperparameters to train all model variations. The optimal hyperparameters are determined using a grid search specific for each task, which are detailed in the Appendix. All model variants are trained on one to four H100 GPUs, depending on task complexity.
@@ 4 Results
Table 1: MixiT performance on algorithmic and sentiment classification tasks, many with a large memorization focus. As shown, both Frozen-QK and MixiT are able solve such tasks. They are competitive with, and can even be superior than, the standard transformer.
Frozen-QK and Frozen-MLP can solve induction heads tasks such as retrieval and hopk.
On the other side of the memorization-heavy tasks on the task spectrum, we test the model variations on tasks that require forming induction heads [17]: needle-in-a-haystack retrieval [37] and $k$-hop induction heads [42]. As shown in Table 2, Frozen-QK is able to solve the tasks on par with the standard fully trained transformer model, reinforcing the observation that induction heads can form even with frozen query and key projectors. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization.
More detailed results on the retrieval task are shown in Figure 2, where we test the performance of each architecture at varying task complexities, controlled by the maximum number of key-value pairs $m_{\max}$ in the sequence. We search for the optimal learning rate for each architecture. To achieve a fair comparison across architectures, we fix the embedding dimension at $n=1024$, with $4$ heads and $2$ layers, which suffice to learn induction heads [17, 16]. As shown, Frozen-QK, Frozen-MLP, and the standard transformer all perform notably better than MixiT as the retrieval complexity increases; and Frozen-QK eventually deteriorates in quality faster than Frozen-MLP and the standard transformer, underscoring the key role attention plays in forming induction heads.
Task \ Model Standard Frozen-MLP Frozen-QK MixiT
$k$-hop Induction Heads $\uparrow$ 99.99% 99.89% 96.73% 48.58%
Retrieval $\uparrow$ 100% 100% 97.01% 11.24%
Table 2: Accuracies for tasks that require forming induction heads. Frozen-QK is able to solve the tasks on par with the standard fully trained transformer model, reinforcing the observation that induction heads can form even with frozen query and key projectors, e.g. as evident in Figure 3. Furthermore, as the attention module plays the key role in forming induction heads [17, 12], Frozen-MLP, with its trainable attention modules, is able to perform to almost perfect accuracy. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization. The retrieval results are for sequences of maximum number of key-value pairs $m_{\text{max}}=30$.
Figure 2: Retrieval accuracy as a function of the number of pairs in the sequence $m_{\max}$, which encodes task complexity. Having trainable queries and keys is necessary to solve the task (Standard and Frozen-MLP). Frozen-QK cannot solve the task for large $m_{\max}$, while MixiT has deteriorating performances.
Frozen-QK can perform competitively on language modeling.
Surprisingly, Frozen-QK comes close to the standard transformer in terms of perplexity, as shown in Table 3. This indicates that trainable attention weights are not always required for successful language modeling. Indeed, as Figure 3 shows, specialized circuits such as induction heads can form even in Frozen-QK. Less surprisingly, MixiT lags behind the standard Transformer, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for language modeling, corroborating earlier works [37, 12].
Model \ Task Wikitext $\downarrow$ Fineweb-edu $\downarrow$
Standard 2.78 3.05
Frozen-QK 3.07 3.16
MixiT 3.73 4.08
Table 3: Performance on language modeling tasks, in terms of log perplexity. Frozen-QK comes surprisingly close in performance to the standard Transformer, despite having random static query and attention weights. MixiT lags behind the standard Transformer, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for good language modeling.
@@
We train an $L=2$, $n_{h}=4$ heads model on the memorization tasks for $10,000$ steps with a learning rate of $0.005$, for a consistent comparison across models. In Table 4 we report the accuracy and the storage capacity bits per parameters. We find that the standard Transformer stores $2.98$ bits per parameters, which is slightly higher than in previous works [1, 59]. Most of the drop occurs in the Frozen-QK model, with $1.13$ bits per parameters, while Frozen-MLP and MixiT have similar storage capabilities at $2.25$ and $2.18$, respectively. Note that these results are yielded in a setting where the accuracies are not saturated at 100%, to give an accurate representation of bits per parameter, rather than a lower bound. Hence they do not contradict the results in Table 1.
@@
Table 4: Standard Transformers outperform all the alternatives in terms of memorization capability, which suggests that MLPs and attention collaborate to remember knowledge. This provides further evidence for recent findings such as knowledge circuits [54] and query localization [10]. Freezing the MLPs causes the most performance drop, indicating that they are the biggest factor when it comes to memorization. Notice that MixiT has slightly more parameters than Frozen-QK because of additional trainable positional embeddings.
These results suggest that (1) the MLPs are largely responsible for memorization, however (2) there is a non-negligible additional contribution given by the integration of MLPs with learnable attention weights. This non-negligible additional contribution provides further evidence for recent findings such as knowledge circuits [54] and query localization [10], in that MLPs and attention collaborate to remember knowledge. In particular, the disproportionately large increase in the learned bits per parameter from Frozen-QK to the fully trained transformer, from 2.25 to 2.98, suggests that the gain in accuracy is more than what can be accounted for by a mere increase in learnable parameters.

### v3

# Is Random Attention Sufficient for Sequence Modeling? Disentangling Trainable Components in the Transformer

@@ Abstract
The transformer architecture is central to the success of modern Large Language Models (LLMs), in part due to its surprising ability to perform a wide range of tasks – including mathematical reasoning, memorization, and retrieval – using only gradient-based learning on next-token prediction. While the core component of a transformer is the self-attention mechanism, we question how much, and which aspects, of the performance gains can be attributed to it. To this end, we compare standard transformers to variants in which either the MLP layers or the attention weights are frozen at initialization. Surprisingly, we find that attention with frozen key and query weights is not only able to form induction heads, but can also perform competitively on language modeling. We formalize this by proving a new expressivity result for transformer models with frozen key and query weights. To further isolate the contribution of attention, we design MixiT, an architecture with entirely random attention scores, with provably stable signal propagation that overcomes prior depth-wise scaling challenges in random transformers. We use the successes and failures of MixiT to understand the role each transformer component plays, such as attention being largely responsible for in-context reasoning, and MLPs being responsible for, but collaborates with attention, on knowledge memorization. Our results suggest that the transformer architecture has a built-in inductive bias towards forming specialized circuits, as it does even without learnable attention weights. [^1] Our code is publicly available at https://github.com/princeton-pli/MixiT.
@@ 1 Introduction
Transformers Vaswani et al. (2017) have rapidly become the workhorse architecture in modern machine-learning systems, powering state-of-the-art models in language, vision, and scientific domains Dosovitskiy et al. (2020); Team et al. (2023); Guo et al. (2025). Their success is typically attributed to the self-attention mechanism, which allows every token to aggregate information from the entire sequence and has been linked to emergent abilities such as long-range retrieval, algorithmic reasoning, and in-context learning. Yet we lack a precise answer to a fundamental question: which degrees of freedom inside the Transformer are truly necessary for these behaviours, and which can be simplified away without harming performance?
Prior work has probed the internals of trained Transformers. Studies of attention maps consistently report the emergence of induction heads that copy information forward and enable in-context retrieval Olsson et al. (2022). Another line of work has focused on whether and how Transformers’ capabilities can emerge Allen-Zhu and Li (2024); Ramesh et al. (2023); Jain et al. (2023) by studying synthetic datasets with Zhong and Andreas (2024) showing that models with frozen self-attention but trainable embeddings still solve many algorithmic tasks. These results hint that different parts of the architecture are responsible for modeling different tasks.
A standard Transformer block, however, contains several interacting components: an attention block, with learnable, input-dependent queries and keys, and MLP blocks composed of fully connected layers. To exemplify the complexity of the interaction, notice that the attention weights can change their value both through changes in the queries/keys parameters, as well as as through the residual-stream representations that feed those projections. Hence, because all the components are trained together, it is hard to discern the contribution of each individual one towards solving a given task. In this work, we study the role of each Transformer’s component by freezing different parts of the architecture to their values at initialization. In partiucular, we consider the following simplified variants:
@@
  • MixiT (Mixing Transformer), a Transformer variant in which the attention sub-layer is replaced by a fixed, randomly initialized mixing matrix. After initialisation the mixing weights are frozen, preventing the model from adapting interactions during training or conditioning them on input content. Learning therefore takes place exclusively in the embedding and MLP blocks and in residual branches.
@@
We compare across several categories of tasks: mathematical reasoning, sentiment classification, memorization, information retrieval, hopk, and language modeling. This allows us to inspect whether and how the different parts of the architecture are able to solve primitive tasks that relate to a range of basic reasoning and memorization skills. Here we summarize our main contributions:
  • We present the surprising finding that Frozen-QK, which has random attention weights, can perform competitively with the standard transformer on language modeling tasks. Indeed, Frozen-QK develops the ability to form specialized circuits such as induction heads during training. This suggests that learnable attention weights are not required to form specialized circuits.
  • We analyze Frozen-QK’s expressiveness, and show that it can indeed learn a wide class of sequence-level functions in Theorem 5.1.
  • We identify an explanation of the subspace selection hypothesis for random transformers (Zhong and Andreas, 2024), namely, representation collapse, and connect it with the phenomenon that random models with standard initializations suffer degenerating performance with respect to depth, through an analysis of the covariance of hidden representations.
  • To resolve this obstruction towards scaling up model depth, we introduce a principled architecture – MixiT, as part of our spectrum of models, and use its successes at pure memorization tasks v.s. failures at pure in-context reasoning tasks such as k-hop, to derive conclusions on the roles of MLPs vs attention. Furthermore, we prove MixiT’s training stability, by analyzing its covariance SDE, in Theorem 2.1.
@@
Our results indicate that the Transformer architecture has some built-in inductive bias towards algorithmic abilities, namely the ability to form specialized circuits, as evidenced by the presence of induction heads even without learnable attention weights. More broadly, our results underpin the importance of having heterogeneous components that are more specialized in efficiently solving modular subtasks, enabling the model to express complex, wide-ranging behaviors.
@@ 2 Models Specification
We consider a decoder-only Transformer based on the widely used Llama architecture Touvron et al. (2023). Given an input sequence $x\in\mathbb{R}^{V\times m}$, where $V$ is the vocabulary size and $m$ is the sequence length, embedded with a linear map to the hidden states $h_{0}=W_{emb}\,x$ where $W_{emb}\in\mathbb{R}^{V\times n}$, where $n$ is the width of the model. At its core, the transformer uses $L$ stacked modules alternating the causal multi-head self-attention layers and MLP layers. Each self attention head is defined as:
@@
where $W_{\ell}^{G},W_{\ell}^{U}\in\mathbb{R}^{{n_{m}\times n}}$ and $W^{D}_{\ell}\in\mathbb{R}^{{n\times n_{m}}}$. $n_{h}$ and $n_{m}$ are the dimensions of the queries/keys for each head and MLP projections. We also use skip connections in both blocks with a pre-normalization scheme Xiong et al. (2020) and causal masking. We apply rotary embeddings Su et al. (2024) to the queries and keys of each layer. As in common practice, we use $n_{m}=4n$ and $Hn_{h}=n$.
@@
We also design a model where the attention map itself is frozen and, to achieve that, input-independent. In the simplest case, this can be obtained by having a random matrix $M_{\ell}\in\mathbb{R}^{m\times m}$ entries with $\mathcal{N}(0,1/m)$ entries, where the factor of $1/m$ acts as a variance-preserving normalizer. To ensure a stable forward pass in terms of depth and width scaling, we follow the principles of attention shaping Noci et al. (2023) and propose the following:
@@
where $W^{M}_{\ell}$ is frozen at initialization and $\bar{W}^{M}_{\ell}$ contains the column-wise empirical average of $W^{M}_{\ell}$, to ensure that each row sums up to 1. In the Appendix, , we show that this attention has a stable forward pass, in the sense that the kernel of the activations has a well-defined depth-and-width limit, converging to a stochastic differential equation (SDE) Li et al. (2022); Noci et al. (2023). When we adopt this architecture, all the weights, excluding the random attention matrix, are trainable. Notably, the following convergence result implies the stability of the forward pass, in particular ruling out the numerical degeneracy such as rank collapse and vanishing gradients (Dong et al., 2021; Noci et al., 2022).
@@
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation˜2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
@@
As the random attention matrix $I+\frac{1}{\sqrt{mn}}W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT does not depend on the input, the rotary positional embedding (Su et al., 2024) standard in Llama models cannot be applied to MixiT, as rotary embeddings are added to learned key and query embeddings, which are not present in MixiT. Hence, we implement a learnable positional embedding for each token position in the sequence, and add it to the corresponding token embedding in the first layer.
@@ 3.1 Tasks
We benchmark on a variety of tasks, covering categories including basic mathematical reasoning, memorization, sentiment classification, and language modeling. The basic mathematical reasoning and memorization tasks are based on tasks used in (Zhong and Andreas, 2024), with increased difficulty on some tasks to better reflect differences between architectures. More experimental details on these tasks can be found in Appendix C.
@@
Each training instance encodes a small randomly generated sequence of pairs followed by a single query, and the model is required to emit the value associated with that query one step later. We uniformly sample a sequence length $m\sim\mathcal{U}\{1,\dots,m_{\max}\}$, where $m_{\max}$ is the maximum sequence length. We then sample $m$ keys $\{k_{i}\}_{i=1}^{m}$ iid from the set $\{\frac{V}{2},\dots,V\}$ and $m$ values $\{v_{i}\}_{i=1}^{m}$ from $\{1,\dots,\tfrac{\text{V}}{2}-1\}$, where $V=256$ is the vocabulary size. The resulting keys are interleaved with their values to form the prefix $(k_{1},\,v_{1}),\,(k_{2},\,v_{2}),\,\dots,\,(k_{m},\,v_{m})]$. A query key $k_{q}$ with $q\in[m]$ is chosen uniformly at random from the keys, and appended to the sequence. The goal is to predict the value corresponding to the query token. This task isolates retrieval ability, and probes associative recall. We sample 40,000 sequences for training, and 4,000 for testing. Transformer-based models typically solve this task by forming induction heads (Olsson et al., 2022).
@@
Following (Sanford et al., 2024), the $k$-hop induction heads task, or hopk, recursively completes bigrams auto-regressively, by repeatedly predicting the token that followed the last occurrence of the currently-considered token. As an example: given the input $X=a{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}dc}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}ad}a$, the 2-hop induction heads prediction is $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c$}. (Sanford et al., 2024) showed that hopk is solvable by a $O(\log k)$-depth transformer. Hence to achieve fair comparison across architectures, we fix the model depth at 5 layers, and search over other hyperparameters.
@@
The parentheses balancing task learns to predict whether a given parentheses sequence is balanced. Within any prefix in the sequence, the number of closing parentheses is less than or equal to the number opening parentheses. Hence this task is solvable by a linear time algorithm. We randomly sample 100,000 parentheses sequences of length 40 for training, while the test set consists of 4,000 such sequences not in the training set. An example is “(()" $\to\text{False}$.
@@
We follow the procedure in Zhong and Andreas (2024): we sample $512^{2}$ key-value pairs, where each key is independently sampled from its value, which is an integer in $[512]$. Because the key-value mapping is random, any success reflects the model’s ability to memorise arbitrary associations. We measure success with the number of bits per parameter that the model can store, defined as $\text{total\_bits}\times\text{model\_acc}/\text{total\_trainable\_params}$, where model_acc is the training accuracy on the task. Notice that for this problem, storing one pattern requires $\log_{2}512=9$ bits, thus $\text{total\_bits}=9\cdot 512^{2}$.
@@
We use the Yelp polarity reviews dataset (Zhang et al., 2015) to test each model variation’s ability to predict a review’s sentiment, i.e. whether a review is positive or negative.
@@
To test the model’s ability to model natural language, we train the model to perform next-token-prediction. We use two datasets, Wikitext-103 (Merity et al., 2016) and Fineweb-edu (Penedo et al., 2024). Wikitext-103 consists of 1,801,350 cleaned-up Wikipedia articles, with test set of size 4358. And Fineweb-edu consists of top-quality entries collected from web-crawled data, focusing on educational text. We randomly sample 1,048,576 entries for training, and 4096 to test.
@@ 3.2 Model training
For each task, we perform a grid search over a range of hyperparameters to train all model variations. The optimal hyperparameters are determined using a grid search specific for each task, which are detailed in the Appendix C. All model variants are trained on one to four H100 GPUs, depending on task complexity.
@@ 4 Results
Frozen-QK can solve induction heads tasks such as retrieval and hopk.
We test our spectrum of models on tasks that require forming induction heads (Elhage et al., 2021): needle-in-a-haystack retrieval (Olsson et al., 2022) and $k$-hop induction heads (hopk) (Sanford et al., 2024). As shown in Table 1, Frozen-QK is able to solve these tasks on par with the standard fully trained transformer model, indicating that induction heads can form even with frozen query and key weights. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization.
Task \ Model Standard Frozen-MLP Frozen-QK MixiT
$k$-hop Induction Heads $\uparrow$ 99.99% 99.89% 96.73% 48.58%
Retrieval $\uparrow$ 100% 100% 97.01% 11.24%
Table 1: Accuracies for tasks that require forming induction heads. Frozen-QK is able to solve the tasks on par with the standard fully trained transformer model, reinforcing the observation that induction heads can form even with frozen query and key projectors, e.g. as evident in Figure 3. Furthermore, as the attention module plays the key role in forming induction heads (Elhage et al., 2021; Crosbie and Shutova, 2025), Frozen-MLP, with its trainable attention modules, is able to perform to almost perfect accuracy. On the other hand, MixiT is unable to solve these tasks, due to its inability to form specialized circuits as the attention scores are frozen at random initialization. The retrieval results are for sequences of maximum number of key-value pairs $m_{\text{max}}=30$.
More detailed results on the retrieval task are shown in Figure˜2, where we test the performance of each architecture at varying task complexities, controlled by the maximum number of key-value pairs $m_{\max}$ in the sequence. As shown, Frozen-QK, Frozen-MLP, and the standard transformer all perform notably better than MixiT as the retrieval complexity increases; and Frozen-QK eventually deteriorates in quality faster than Frozen-MLP and the standard transformer, underscoring the key role attention plays in forming induction heads.
Figure 2: Retrieval accuracy as a function of the number of pairs in the sequence $m_{\max}$, which encodes task complexity. While MixiT has rapidly deteriorating performance with respect to task complexity, Frozen-QK, with its ability to form induction heads, reaches its performance ceiling much more slowly.
Frozen-QK can perform competitively on language modeling.
Surprisingly, Frozen-QK comes close to the standard transformer in terms of perplexity, as shown in Table 2. This indicates that trainable attention weights are not always required for successful language modeling. Indeed, as Figure 3 shows, specialized circuits such as induction heads can form even in Frozen-QK. Less surprisingly, MixiT lags behind the standard Transformer, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for language modeling, corroborating earlier works (Olsson et al., 2022; Crosbie and Shutova, 2025).
Model \ Task Wikitext $\downarrow$ Fineweb-edu $\downarrow$
Standard 2.78 3.05
Frozen-QK 3.07 3.16
MixiT 3.73 4.08
Table 2: Performance on language modeling tasks, in terms of log perplexity. Frozen-QK comes surprisingly close in performance to the standard Transformer, despite having random static query and attention weights. MixiT has notably worse performance, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for good language modeling.
@@
Table 3: MixiT performance on algorithmic and sentiment classification tasks. As shown, both Frozen-QK and MixiT are able solve such tasks. They are competitive with, and can even be superior than, the standard transformer.
@@
Table˜4 shows the accuracy and the storage capacity via bits per parameter. We find that the standard transformer stores $2.98$ bits per parameters, which is slightly higher than in previous works Allen-Zhu and Li (2024); Zhong and Andreas (2024). Most of the drop occurs in the Frozen-QK model, with $1.13$ bits per parameters, while Frozen-MLP and MixiT have similar storage capabilities at $2.25$ and $2.18$, respectively. Note that these results are yielded in a setting where the accuracies are not saturated at 100%, to give an accurate representation of bits per parameter. Hence they do not contradict the results in Table 3.
@@
Table 4: Standard Transformers outperform all the alternatives in terms of memorization capability, which suggests that MLPs and attention collaborate to remember knowledge. This provides further evidence for recent findings such as knowledge circuits (Yao et al., 2024) and query localization (Chen et al., 2025). Freezing the MLPs causes the most performance drop, indicating that they are the biggest factor when it comes to memorization. Notice that MixiT has slightly more parameters than Frozen-QK because of additional trainable positional embeddings.
These results suggest that (1) the MLPs are largely responsible for memorization, however (2) there is a non-negligible additional contribution given by the integration of MLPs with learnable attention weights. This non-negligible additional contribution provides further evidence for recent findings such as knowledge circuits (Yao et al., 2024) and query localization (Chen et al., 2025), in that MLPs and attention collaborate to remember knowledge. In particular, the disproportionately large increase in the learned bits per parameter from Frozen-QK to the fully trained transformer, from 2.25 to 2.98, suggests that the gain in accuracy is more than what can be accounted for by a mere increase in learnable parameters.